# DN-Splatter: Depth and Normal Priors for Gaussian Splatting and Meshing

### v2

@@ 5.3 Ablation studies
Multi-view depth [59] .3120 .2375 .2748 .6408 .1903
Patch-based depth [52] .1183 .1766 .1474 .7975 .2236
@@
Table 6: Ablation of depth supervision. We compare monocular [3, 15], multi-view [15], relative (patch-based) [52], and sensor depth supervision strategies on the ”b20a261fdf” scene of ScanNet++.
Despite the low-resolution sensor depths lacking detail and containing inaccuracies (mostly at object edges), they remain practical for use in real-world indoor scenes. Future research is needed to bridge the performance gap for monocular depth supervision.
Mesh extraction techniques. Lastly, we investigate various Poisson meshing techniques. In Table 7, we demonstrate that extracting oriented point sets from optimized depth and normal maps results in smoother and more realistic reconstructions compared to other methods. We report mesh evaluation metrics for these different techniques. We compare several approaches: directly using trained Gaussian means and normals (total of 512k Gaussians); extraction of surface density at levels 0.1 and 0.5, as proposed in SuGaR  [12]; back-projection of optimized depth and normal maps, as explained in Section 4.4. All models were trained with our depth and normal regularization. To ensure a fair comparison, we set the total number of extracted points to 500k for both the surface density and back-projection methods.
(a) GaussiansRefer to caption (b) Density 0.1Refer to caption (c) Density 0.5Refer to caption (d) OursRefer to caption (e) GTRefer to caption
Acc. $\downarrow$ Comp. $\downarrow$ C-$L_{1}$downarrow NC $\uparrow$ F-score $\uparrow$
Gaussians .0206 .0412 .0309 .9091 .9117
SuGaR [12]: density 0.1 .0130 .0357 .0243 .9301 .9275
SuGaR [12]: density 0.5 .0083 .0304 .0193 .9309 .9325
Back-projection (ours) .0074 .0312 .0194 .9428 .9310
Table 7: Ablation of Poisson mesh extraction techniques: Replica. We compare naive Gaussian-based meshing, the meshing strategy from [12], and our back-projection approach. All models were trained using the proposed depth and normal objectives.
@@ 6 Conclusion
We presented DN-Splatter, a method for depth and normal regularization of 3D Gaussian splatting. This simple yet effective strategy enhances photorealism by improving common novel view synthesis metrics and significantly improving depth estimation and surface quality extracted from the Gaussian scene. We demonstrated that prior regularization is essential for achieving more geometrically valid and consistent reconstructions in challenging indoor scenes. Although we improved over the state-of-the-art methods, our focus was limited to densely captured scenes with relatively still cameras. Future work could address more challenging and sparser data captures, which often suffer from motion blur and other capture artifacts. Additionally, better meshing techniques are needed to optimize Gaussian scene parameters and mesh quality concurrently, as Poisson surface reconstruction is more sensitive to the estimated positions of the points than to their normal estimates. Furthermore, more research is needed to bridge the performance gap between monocular depth estimation and real sensor depth supervision. This remains an area for future work.
@@
[…]
1 ETH Zurich, 2 Tampere University, 3 Aalto University, 4 Spectacular AI
Corresponding author: matiasturkulainen@gmail.com
@@ A.2 Datasets
ScanNet++. We use the ”b20a261fdf” and ”8b5caf3398” scenes in our experiments. We use the iPhone sequences with COLMAP registered poses. The sequences contain 358 and 705 registered images respectively. We uniformly load every 5th frame from the sequences from which we reserve every 10th frame for evaluation.
@@ B.3 Depth losses
(a) We load every 3/5/8/12 views from the whole training sequence (around 260). Results are evaluated on ”Courtroom” from Tanks & Temples.
@@
(b) We load every 5/8/12/20 views from the whole training sequence (around 270). Results are evaluated on ”8b5caf3398” from ScanNet++ DSLR sequence.
@@
Table 12: Comparison of DN-Splatter performance with monocular depth supervision. We ablate the Zoe-Depth[3] and DepthAnything[53] monocular estimators with sparse views on the ”Courtroom” sequence of Tanks & Temples advanced dataset. Monocular depth supervision aids in novel-view synthesis under sparse settings.
Appendix C Additional quantitative results
Here we provide additional quantitative results for DN-Splatter. In Table 11, we show the depth evaluation performance of our proposed regularization scheme on the MuSHRoom dataset, evaluated against ground truth Faro lidar scanner data instead of the low-resolution iPhone depths. This corresponds to Table 3 from the main paper, which compares depth metrics on iPhone depth captures for the same scenes and baselines. When comparing to laser scanner depths, our method still out performs other baseline methods on depth estimation.
We also consider the performance of DN-Splatter supervised by only monocular depth estimates on a large scale Tanks & Temples scene in Table 12. We consider training with dense and sparse captures and conclude that although monocular depth supervision in dense captures provides minimal improvements, the increase in novel view synthesis under sparse settings is notable.
Lastly, in Table 13 we compare the performance of various depth losses described in Section B.3 on depth estimation and novel view synthesis. There are several interesting observations. First, the logarithmic depth loss $L_{\text{LogL1}}$ outperforms other popular variants like $L_{\text{L1}}$ or $L_{\text{MSE}}$ on depth and RGB synthesis. Second, the gradient-aware logarithmic depth variant $L_{\hat{D}}$ outperforms the simpler variant, validating our assumption that captured sensor depths, like those from iPhone cameras, tend to contain noise and inaccuracies at edges or sharp boundaries. Therefore, the gradient-aware variant mitigates these inaccurate sensor readings.
@@ Appendix D Additional qualitative results
Figure 7: Qualitative comparison of rendered depth and RGB images. Comparison of baseline methods on the ”sauna” sequence from the MuSHRoom dataset.
@@
Figure 8: Qualitative comparison of rendered depth and RGB images. Comparison of baseline methods on the ”classroom” and ”coffee room” sequences from the MuSHRoom dataset.
@@
Figure 9: Qualitative comparison of rendered depth and RGB images. Comparison of baseline methods on the ”koivu” sequence from the MuSHRoom dataset.

### v3

@@ 5.3 Ablation studies
Multi-view Stereo (MVSNet) [60] .3120 .2375 .2748 .6408 .1903
Pearson Loss  [52] w/ Metric3D .1183 .1766 .1474 .7975 .2236
@@
Table 8: Alternative depth regularization strategies. We compare regularization with monocular [3, 15] and Multi-view Stereo depth estimates  [15] using our $\mathcal{L}_{\hat{D}}$ loss, the Pearson correlation loss (patch-based) [52] for relative depth supervision with Zoe-Depth estimates, as well as utilizing iPhone sensor depth supervision on the ’b20a261fdf’ scene from ScanNet++.
@@ 6 Conclusion
We presented DN-Splatter, a method for depth and normal regularization of 3DGS to address photorealistic and geometrically accurate reconstruction of challenging indoor datasets. This simple yet effective strategy enhances novel-view metrics and significantly improves the surface quality extracted from a Gaussian scene. We demonstrated that prior regularization is essential for achieving more geometrically valid and consistent reconstructions in challenging indoor scenes.
7 Acknowledgments
We thank Tobias Fischer, Songyou Peng, and Philipp Lindenberger for their fruitful discussions. We thank Jiepeng Wang and Marcus Klasson for their help in proof reading. We acknowledge funding from the Academy of Finland (grant No. 327911, 353138, 324346, and 353139) and support from the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Allice Wallenberg Foundation. MT was funded by the Finnish Center for Artificial Intelligence (FCAI).
@@ A.1 Baselines
In addition, we have modified the original source-code to support depth rendering, which was not possible in the original author’s code release. This is achieved by replacing the CUDA backend with a variant that also includes depth rendering support.
@@
2DGS + $\mathcal{L_{\hat{D}}}$ variant. We implement the proposed depth regularization strategy into the official 2DGS code release. Specifically, we enable supervision and gradient flow to depths within the CUDA backend rasterizer and supervise with sensor or monocular depth estimates. Our overall optimization loss becomes $\mathcal{L}=\mathcal{L}_{\text{rgb}}+\lambda_{d}\mathcal{L}_{d}$ where $\lambda_{d}$ is set to 0.2 and $\mathcal{L}_{\text{rgb}}$ is the original loss from [16].
@@ A.2 Datasets
ScanNet++. We use the "b20a261fdf" and "8b5caf3398" scenes in our experiments. We use the iPhone sequences with COLMAP registered poses. The sequences contain 358 and 705 registered images respectively. We uniformly load every 5th frame from the sequences from which we reserve every 10th frame for evaluation.
@@ B.3 Depth losses
Appendix C Additional quantitative results
Here we provide additional quantitative results for DN-Splatter. We provide a comparison of Poisson meshing strategies, comparison of depth estimation quality with ground truth Faro scanner data, as well as further ablations on depth loss variants.
C.1 Mesh extraction techniques
We investigate various Poisson meshing techniques. In Table 12, we demonstrate that extracting oriented point sets from optimized depth and normal maps results in smoother and more realistic reconstructions compared to other methods. We report mesh evaluation metrics for these different techniques. We compare several approaches: directly using trained Gaussian means and normals for Poisson meshing (total of 512k Gaussians); extraction of surface density at levels 0.1 and 0.5 by projecting rays from camera views and querying scene intersections based on local density values, as proposed in SuGaR  [11]; and back-projection of optimized depth and normal maps. All models were trained with our depth and normal regularization. To ensure a fair comparison, we set the total number of extracted points to 500k for both the surface density and back-projection methods.
(a) Gaussians[Uncaptioned image] (b) Density 0.1[Uncaptioned image] (c) Density 0.5[Uncaptioned image] (d) Ours[Uncaptioned image] (e) GT[Uncaptioned image]
Acc. $\downarrow$ Comp. $\downarrow$ C-$L_{1}$downarrow NC $\uparrow$ F-score $\uparrow$ Gaussians .0206 .0412 .0309 .9091 .9117 SuGaR [11]: density 0.1 .0130 .0357 .0243 .9301 .9275 SuGaR [11]: density 0.5 .0083 .0304 .0193 .9309 .9325 Back-projection (ours) .0074 .0312 .0194 .9428 .9310
Table 12: Ablation of Poisson mesh extraction techniques: Replica. We compare naive Gaussian-based meshing, the meshing strategy proposed in SuGaR [11], and our back-projection approach. All models were trained using the proposed depth and normal objectives.
@@ C.1 Mesh extraction techniques
(a) We load every 3/5/8/12 views from the whole training sequence (around 260). Results are evaluated on "Courtroom" from Tanks & Temples.
@@
(b) We load every 5/8/12/20 views from the whole training sequence (around 270). Results are evaluated on "8b5caf3398" from ScanNet++ DSLR sequence.
@@
Table 14: Comparison of DN-Splatter performance with monocular depth supervision. We ablate the Zoe-Depth[3] and DepthAnything[53] monocular estimators with sparse views on the "Courtroom" sequence of Tanks & Temples advanced dataset. Monocular depth supervision aids in novel-view synthesis under sparse settings.
C.2 Depth estimation compared to Faro scanner ground truth
In Table 13, we show the depth evaluation performance of our proposed regularization scheme on the MuSHRoom dataset, evaluated against ground truth Faro lidar scanner data instead of the low-resolution iPhone depths. This corresponds to Table 3 from the main paper, which compares depth metrics on iPhone depth captures for the same scenes and baselines. When comparing to laser scanner depths, our method still out performs other baseline methods on depth estimation.
C.3 Additional depth comparisons
We consider the performance of DN-Splatter within sparse view setting guided by only monocular depth estimates. We test on the large scale Tanks & Temples scene in Table 14. We consider training with dense and sparse captures and conclude that although monocular depth supervision in dense captures provides minimal improvements, the increase in novel view synthesis metrics under sparse view settings is notable. Lastly, in Table 15 we compare the performance of various depth losses described in Section B.3 on depth estimation and novel view synthesis. There are several interesting observations. First, the logarithmic depth loss $L_{\text{LogL1}}$ outperforms other popular variants like $L_{\text{L1}}$ or $L_{\text{MSE}}$ on depth and RGB synthesis. Second, the gradient-aware logarithmic depth variant $L_{\hat{D}}$ outperforms the simpler variant, validating our assumption that captured sensor depths, like those from iPhone cameras, tend to contain noise and inaccuracies at edges or sharp boundaries. Therefore, the gradient-aware variant mitigates these inaccurate sensor readings.
@@ Appendix D Additional qualitative results
D.1 Normal smoothing loss
We visualize the impact of $\mathcal{L}_{\text{smooth}}$ prior on rendered normal estimates in Fig. 7. We achieve smoother predictions with the prior.
(a) RGB
(b) w/o $\mathcal{L}_{\text{smooth}}$
(c) w/ $\mathcal{L}_{\text{smooth}}$
(d) Omnidata GT
Figure 7: Qualitative comparison of normal smoothing prior. We visualize normal estimates with and without our $\mathcal{L}_{\text{smooth}}$ smoothing prior on the ’VR Room’ scene from MuSHRoom dataset.
D.2 2DGS vs. DN-Splatter renders
In Fig. 8 we compare novel-view and depth estimation renders using baseline Splatfacto and 2DGS models as well as a variant of 2DGS with depth supervision enabled and our method.
(a) iPhone RGB
(b) Splatfacto
(c) 2DGS
(d) 2DGS w/ Sensor Depth
(e) Ours w/ Sensor Depth
(f) iPhone Depth
(g) Splatfacto
(h) 2DGS
(i) 2DGS w/ Sensor Depth
(j) Ours w/ Sensor Depth
(k) iPhone RGB
(l) Splatfacto
(m) 2DGS
(n) 2DGS w/ Sensor Depth
(o) Ours w/ Sensor Depth
(p) iPhone Depth
(q) Splatfacto
(r) 2DGS
(s) 2DGS w/ Sensor Depth
(t) Ours w/ Sensor Depth
Figure 8: Qualitative comparison between 2DGS and DN-Splatter. We supervise both 2DGS and our method with the $\mathcal{L}_{\hat{D}}$ regularization loss and visualize novel-view and depth renders from the ’Honka’ and ’Kokko’ scenes from the MuSHRoom dataset. We note that DN-Splatter obtains higher metrics in both novel-view and mesh reconstruction metrics; whilst 2DGS obtains more smooth depth renders.
D.3 Mesh and NVS renders
@@
Figure 10: Qualitative comparison of rendered depth and RGB images. Comparison of baseline methods on the "sauna" sequence from the MuSHRoom dataset.
@@
Figure 11: Qualitative comparison of rendered depth and RGB images. Comparison of baseline methods on the "classroom" and "coffee room" sequences from the MuSHRoom dataset.
@@
Figure 12: Qualitative comparison of rendered depth and RGB images. Comparison of baseline methods on the "koivu" sequence from the MuSHRoom dataset.